# UniGS: Unified Geometry-Aware Gaussian Splatting for Multimodal Rendering

### v1

@@ 4.3 Semantic Rendering Evalution
We compare our method with the SGS-SLAM [32] framework, which employs a task-specific Gaussian map representation and is supervised using semantic result images for semantic tasks, differing from our unified framework that predicts semantic logits. The comparison across multiple scenes is shown in Fig. 5 and Tab. 3. From the quantitative results, it can be observed that our method achieves superior performance in both image rendering quality (PSNR) and semantic segmentation quality (mIoU) within a very short training time. Thanks to our unified rendering architecture, the image rendering FPS is also fast. Additional results can be found in our supplementary video.
@@
GeoGaussian [33] 126021 440.9 94755 441.1 53457 436.2 100877 423.6
3DGS [25] 504167 128.7 421167 136.9 285394 162.4 543581 137.5
Scaffold-GS [40] 301839 240.1 247081 249.1 182304 324.0 235446 346.2
Normal-GS [61] 318064 180.3 226283 178.7 185243 250.7 204358 180.0
Ours 161081 169.8 119947 163.1 48970 296.7 169273 197.8
@@ 4.4 Runtime and Memory Cost Analysis
In Tab. 4, we present a comparison of the Count of Gaussians and rendering FPS against other methods. Our approach employs an effective differentiable pruning module, which removes a significant number of redundant Gaussians. As a result, the number of Gaussian splats in our reconstructed scenes is the smallest. However, since our method requires rendering multiple modalities simultaneously, the FPS experiences a certain degree of decrease. Nevertheless, this does not severely hinder our rendering speed. In some scenarios, our rendering speed remains superior even compared to Normal-GS [61].
@@ 4.5 Ablation Study
Figure 7: Visualization of the distribution of the gradient factor. To enhance the visualization of the gradient factor, the brightness of all Gaussians is uniformly minimized, resulting in blue regions as rendered by SuperSplat [8]. (a) Scene with original brightness. (b) Areas highlighted in yellow indicate regions where the gradient factor $k>1.5$. (c) Areas highlighted in yellow indicate regions where $k<0.5$.
@@ 6.6 Tile-Based Gradient Backpropagation
In the previous sections, we obtain the losses for the semantic logits, depth, normal, and gradient factor map with respect to their respective ground truths, which are given by $\frac{\partial\mathcal{L}}{\partial\tilde{\mathbf{O}}}$, $\frac{\partial\mathcal{L}}{\partial\tilde{\mathbf{D}}}$, $\frac{\partial\mathcal{L}}{\partial\tilde{\mathbf{N}}}$, and $\frac{\partial\mathcal{L}}{\partial\tilde{\mathbf{K}}}$. Corresponding to Alg. 2, we optimize the Gaussian parameters through the backward propagation of gradients via the alpha rendering process. The backward propagation pipeline is illustrated in Alg. 4. We omit the gradient propagation process for RGB (indicated as ...), which is consistent with the original implementation [25].
@@
iNPG [44] 26.43 0.725 0.339 21.72 0.723 0.330 23.62 0.797 0.423 - - -
Mip-NeRF360 [1] 29.23 0.844 0.207 22.22 0.759 0.257 29.40 0.901 0.245 - - -
RaySplats [3] 27.31 0.846 0.237 22.20 0.829 0.202 29.57 0.900 0.320 - - -
GeoGaussian [33] 26.26 0.813 0.219 20.00 0.802 0.261 26.07 0.873 0.310 24.55 0.946 0.114
3DGS [25] 28.69 0.870 0.182 23.14 0.841 0.183 29.41 0.903 0.243 36.50 0.940 0.120
Scaffold-GS [40] 28.84 0.848 0.220 23.96 0.853 0.177 30.21 0.906 0.254 37.45 0.951 0.090
Normal-GS [61] 29.34 0.869 0.194 24.22 0.854 0.174 30.19 0.910 0.252 38.19 0.977 0.043
Ours 30.10 0.887 0.148 23.57 0.860 0.168 29.66 0.899 0.231 36.68 0.962 0.083
@@ 7.1 RGB Rendering Evaluation
To validate the effectiveness of our method in providing valuable information for NVS, we compare its performance on novel view RGB rendering with state-of-the-art frameworks [1] [44] [73] [25] [40] [33]. All methods utilize their provided original code and unified original data. The qualitative results of the RGB rendering are shown in Fig. 8, and the quantitative comparisons are presented in Tab. 5. It can be observed that our method achieves state-of-the-art results in novel view synthesis for certain scenes. By incorporating geometric constraints into the framework, our algorithm achieves SOTA geometric performance while maintaining RGB reconstruction quality and significantly improving memory and rendering efficiency.
@@ 7.2 Statistical Outlier Analysis
To quantify the quality of predicted point clouds relative to ground truth data in Replica [52], we employ a statistical approach based on Z-score analysis of point-to-point distances. This method allows us to identify and measure outliers in predicted point clouds with respect to the ground truth mesh points.
@@ 7.3 More Results on Geometric Rendering
It is worth noting that in the Mip-NeRF [1], Tanks & Temples [27], and Deep Blending [16] datasets, the depth and normal information used for supervision is provided by Normal-GS [61] and may still be inaccurate in certain scenarios. Nevertheless, our method achieves competitive results in both depth and normal estimation.
@@ 7.4 Ablation on Algorithm Components
Figure 9: Quantitative comparison of depth and normal estimation accuracy across different methods on several datasets.

### v2

@@ 4.3 Semantic Rendering Evalution
We compare our method with the SGS-SLAM [sgsslam] framework, which employs a task-specific Gaussian map representation and is supervised using semantic result images for semantic tasks, differing from our unified framework that predicts semantic logits. The comparison across multiple scenes is shown in Fig. 5 and Tab. 3. From the quantitative results, it can be observed that our method achieves superior performance in both image rendering quality (PSNR) and semantic segmentation quality (mIoU) within a very short training time. Thanks to our unified rendering architecture, the image rendering FPS is also fast. Additional results can be found in our supplementary video.
@@
GeoGaussian [li2024geogaussian] 126021 440.9 94755 441.1 53457 436.2 100877 423.6
3DGS [3dgs] 504167 128.7 421167 136.9 285394 162.4 543581 137.5
[…]
@@ 4.4 Runtime and Memory Cost Analysis
In Tab. 4, we present a comparison of the Count of Gaussians and rendering FPS against other methods. Our approach employs an effective differentiable pruning module, which removes a significant number of redundant Gaussians. As a result, the number of Gaussian splats in our reconstructed scenes is the smallest. However, since our method requires rendering multiple modalities simultaneously, the FPS experiences a certain degree of decrease. Nevertheless, this does not severely hinder our rendering speed. In some scenarios, our rendering speed remains superior even compared to Normal-GS [normalgs].
@@ 4.5 Ablation Study
Figure 7: Visualization of the distribution of the gradient factor. To enhance the visualization of the gradient factor, the brightness of all Gaussians is uniformly minimized, resulting in blue regions as rendered by SuperSplat [supersplat]. (a) Scene with original brightness. (b) Areas highlighted in yellow indicate regions where the gradient factor $k>1.5$. (c) Areas highlighted in yellow indicate regions where $k<0.5$.
@@ 6.6 Tile-Based Gradient Backpropagation
In the previous sections, we obtain the losses for the semantic logits, depth, normal, and gradient factor map with respect to their respective ground truths, which are given by $\frac{\partial\mathcal{L}}{\partial\tilde{\mathbf{O}}}$, $\frac{\partial\mathcal{L}}{\partial\tilde{\mathbf{D}}}$, $\frac{\partial\mathcal{L}}{\partial\tilde{\mathbf{N}}}$, and $\frac{\partial\mathcal{L}}{\partial\tilde{\mathbf{K}}}$. Corresponding to Alg. 2, we optimize the Gaussian parameters through the backward propagation of gradients via the alpha rendering process. The backward propagation pipeline is illustrated in Alg. 4. We omit the gradient propagation process for RGB (indicated as ...), which is consistent with the original implementation [3dgs].
@@
iNPG* [muller2022instant] 26.43 0.725 0.339 21.72 0.723 0.330 23.62 0.797 0.423 - - -
[…]
RaySplats* [raysplat] 27.31 0.846 0.237 22.20 0.829 0.202 29.57 0.900 0.320 - - -
GeoGaussian [li2024geogaussian] 26.26 0.813 0.219 20.00 0.802 0.261 26.07 0.873 0.310 24.55 0.946 0.114
[…]
2DGS [huang20242d] 27.34 0.814 0.234 22.89 0.833 0.206 27.96 0.873 0.281 36.47 0.943 0.107
[…]
@@ 7.1 RGB Rendering Evaluation
To validate the effectiveness of our method in providing valuable information for NVS, we compare its performance on novel view RGB rendering with state-of-the-art frameworks [barron2022mip] [muller2022instant] [yu2021plenoxels] [3dgs] [scaffoldgs] [li2024geogaussian]. All methods utilize their provided original code and unified original data. The qualitative results of the RGB rendering are shown in Fig. 8, and the quantitative comparisons are presented in Tab. 5. It can be observed that our method achieves state-of-the-art results in novel view synthesis for certain scenes. By incorporating geometric constraints into the framework, our algorithm achieves SOTA geometric performance while maintaining RGB reconstruction quality and significantly improving memory and rendering efficiency.
@@ 7.2 Statistical Outlier Analysis
To quantify the quality of predicted point clouds relative to ground truth data in Replica [straub2019replica], we employ a statistical approach based on Z-score analysis of point-to-point distances. This method allows us to identify and measure outliers in predicted point clouds with respect to the ground truth mesh points.
@@ 7.3 More Results on Geometric Rendering
It is worth noting that in the Mip-NeRF [barron2022mip], Tanks & Temples [knapitsch2017tanks], and Deep Blending [hedman2018deep] datasets, the depth and normal information used for supervision is provided by Normal-GS [normalgs] and may still be inaccurate in certain scenarios. Nevertheless, our method achieves competitive results in both depth and normal estimation.
@@ 7.4 Ablation on Algorithm Components
Figure 9: Qualitative comparison of depth and normal estimation accuracy across different methods on several datasets.